OpenStudy (anonymous):

Integrate (x)/(1+x^2) DX ....show work please

OpenStudy (anonymous):

u - sub for this one. put \[u=1+x^2, du=2xdx,\frac{1}{2}du = dx\] and do the rest mostly in your head

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=1%2F2+log%281%2Bx^2%29&lk=1&a=ClashPrefs_*Math- There is full calculation more than i can explain ... =)

OpenStudy (anonymous):

\[\int\frac{du}{u}=\ln(u)\] replace by u by \[x^2+1\] and dont forget the one half out front and +C at the end

OpenStudy (anonymous):

Srry this is the one.. http://www.wolframalpha.com/input/?i=%28x%29%2F%281%2Bx^2%29

OpenStudy (anonymous):

Thanks satellite, I'm still not completely comfortable with the application of U.

OpenStudy (anonymous):

did you get this one?

OpenStudy (anonymous):

Yes I understand it, the only part of integration that becomes confusing for me is selecting which part of the equation to turn into U.

OpenStudy (anonymous):

ah i see. it will make more sense if you take the derivative of your answer and see why it works. "u-sub" just a gimmick for doing the chain rule backwards. so usually you use it when you see a composite function times its derivative.

OpenStudy (anonymous):

hmm, trying to do equation....still a bit lost here.

OpenStudy (anonymous):

if you take the derivative of \[\ln(x^2+1)\] you see that you get \[\frac{1}{x^2+1}\times 2x=\frac{2x}{x^2+1}\] by the chain rule. this is almost exactly what you want, you are just off by a factor of 2

OpenStudy (anonymous):

that is what the "u-sub' allows you to do. see what you need to adjust by to get the right thing

OpenStudy (anonymous):

how did you get to ln(x^2 + 1)

OpenStudy (anonymous):

ok lets to slow

OpenStudy (anonymous):

your job is \[\int\frac{x}{x^2+1}dx\] meaning you are looking for some function whose derivative is \[\frac{x}{x^2+1}\]

OpenStudy (anonymous):

in the denominator you have \[x^2+1\] and in the numerator you have \[x\] which is almost the derivative of the denominator. you are only off by a factor of 2

OpenStudy (anonymous):

this is a set up for a u - substitution, which is device for running the chain rule backwards. so you can say: put \[u=x^2+1\] and then \[du=2xdx\] so \[\frac{1}{2}du=dx\] and now rewrite your integral entirely in terms of u as \[\frac{1}{2}\int\frac{1}{u}du\]

OpenStudy (anonymous):

thats where im confused

OpenStudy (anonymous):

what does this buy you? well now you know exactly an anti-derivative of \[\frac{1}{u}\] namely \[\ln(u)\] because \[\frac{d}{du}\ln(u)=\frac{1}{u}\]

OpenStudy (anonymous):

why isnt it x/u instead of 1/u

OpenStudy (anonymous):

sorry i meant you have written \[2xdx=du\implies xdx=\frac{1}{2}du\]

OpenStudy (anonymous):

ahhhhh, X's cancel out?

OpenStudy (anonymous):

so now you have everything in terms of u. no they don't "cancel" they just disappear. it is a trick, a trick to write your integral in a form you can answer

OpenStudy (anonymous):

again let me stress this is just a gimmick to get what you want. you want a function whose derivative is \[\frac{x}{x^2+1}\] so how are you going to find it? once we recognize that he denominator is almost the derivative of the numerator we know that it is going to be \[\ln(x^2+1)\] but we need to know how to adjust.

OpenStudy (anonymous):

multiply by the 1/2 left over and its your answer right?

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

here is another maybe simpler example. suppose you want \[\int x\cos(x^2)dx\]

OpenStudy (anonymous):

u= x^2?

OpenStudy (anonymous):

so you know that the derivative of sine is cosine and you are almost home free. you know your answer will be \[\sin(x^2)\] basically. but the derivative of \[\sin(x^2)\] is \[2x\cos(x)\] which is not exactly what you have. so you have to divide by two. you are right. you would say \[u=x^2\] \[du=2xcx\] \[\frac{1}{2}du=xdx\] and now you have exactly what you want

OpenStudy (anonymous):

ok thanks, I actually havent seen sin or cos yet in calc, but prob will soon

OpenStudy (anonymous):

oh well good luck! and if you get confused check your answer by differentiation

OpenStudy (anonymous):

thankfully I have the ti 89 to help me out as well haha, thanks